# From Prompts to Printable Models: Support-Effective 3D Generation via Offset Direct Preference Optimization

### v1

@@ Abstract
The transition from digital 3D models to physical objects via 3D printing often requires support structures to prevent overhanging features from collapsing during the fabrication process. While current slicing technologies offer advanced support strategies, they focus on post-processing optimizations rather than addressing the underlying need for support-efficient design during the model generation phase. This paper introduces SEG (Support-Effective Generation), a novel framework that integrates Direct Preference Optimization with an Offset (ODPO) into the 3D generation pipeline to directly optimize models for minimal support material usage. By incorporating support structure simulation into the training process, SEG encourages the generation of geometries that inherently require fewer supports, thus reducing material waste and production time. We demonstrate SEG’s effectiveness through extensive experiments on two benchmark datasets, Thingi10k-Val and GPT-3DP-Val, showing that SEG significantly outperforms baseline models such as TRELLIS, DPO, and DRO in terms of support volume reduction and printability. Qualitative results further reveal that SEG maintains high fidelity to input prompts while minimizing the need for support structures. Our findings highlight the potential of SEG to transform 3D printing by directly optimizing models during the generative process, paving the way for more sustainable and efficient digital fabrication practices.
@@ I Introduction
Imagine a world where the digital creations that spark our imagination can seamlessly transition from the screen to our hands. In recent years, the rapid advancements in 3D generation models, particularly those utilizing diffusion techniques, have opened the doors to unprecedented possibilities in digital fabrication. These models, capable of producing intricate and diverse geometries from simple textual descriptions, are not just tools for artists and designers; they represent a paradigm shift in how we conceive and interact with objects in our environment.
As we increasingly rely on digital assets, the potential to transform these virtual creations into tangible forms becomes a vital step in enhancing our daily lives. Personalized fabrication through 3D printing allows us to materialize ideas that were once confined to screens, creating a bridge between the digital and physical worlds. Whether it’s custom-designed home decor, medical implants tailored to individual patients, or educational tools that bring concepts to life, the implications of this technology are profound. The ability to produce unique, on-demand objects can radically change how we think about consumption, creativity, and accessibility.
However, this exciting potential is accompanied by significant challenges. One of the most pressing issues in translating 3D-generated models into printable objects is the necessity of support structures. Existing text-to-3D models ignore printability constraints, while slicers handle support after geometry is fixed. These supports are crucial for maintaining the integrity of overhanging features during the printing process. Yet, they often lead to material waste and extended production times, and they can compromise the surface quality of the final product [1, 2]. Traditional geometric optimization approaches have proven effective by either deforming the 3D model in the post-processing phase [3] or decomposing the model into a sequence of subparts for multi-axis fabrication [4, 5]. Unfortunately, many existing 3D generation models fail to account for these practical constraints during their training, resulting in geometries that, while visually stunning, may not be feasible for real-world fabrication.
Recognizing these challenges, we set out to create a solution that not only enhances the capabilities of 3D generation but also aligns them with the practical realities of 3D printing. We introduce SEG, a novel framework that employs Direct Preference Optimization with an Offset (ODPO) [6] to refine generative models specifically for support-efficient object creation. By integrating simulation-based feedback on support structure volume into the alignment objective, SEG encourages the generation of geometries that require fewer supports while maintaining the fidelity and aesthetic appeal of the designs. Our contributions can be summarized as follows:
@@ II-A Support-oriented Modeling and Optimization
Prior research in support-aware modeling primarily focused on post-processing strategies that modify a completed model to reduce the need for supports or to enhance its structural stability. For example, build-orientation optimization algorithms evaluate various possible orientations of a given part, seeking to minimize the area of unsupported faces or to reduce the average angle of overhangs [8, 9]. Once an orientation is selected, explicit support-generation methods design and position support structures—ranging from tree-like branches [10] to locally adaptive struts that are strategically placed according to geometric needs [11, 12]. These algorithms aim to uphold problematic features while balancing the trade-offs between print duration and material efficiency. Another class of techniques, partitioning-based methods, decompose complex models into smaller, more manageable sub-parts that can each be printed with fewer supports, subsequently reassembling them via geometric joints or adhesives [13].
By contrast, more recent approaches integrate support constraints directly into the modeling phase, rather than treating support generation as a secondary step. For instance, topology optimization under overhang constraints incorporates additional penalties—such as those based on feature angles or surface curvature—into classical density-based solvers, thereby discouraging the formation of unprintable regions during the design process [14, 15]. Other methods operate directly on the mesh geometry, applying curvature-driven smoothing to eliminate sharp overhangs, or enforcing manufacturability criteria such as minimum wall thickness and maximum unsupported span [3].
Figure 1: Overview of the proposed SEG pipeline. The process begins with data sampling from the Thingi10k database, where multi-view images are generated and captioned using Cap3D. The Text-to-Mesh transformation utilizes a sparse flow transformer $G_{L}$ for model generation. Subsequently, the ODPO training phase incorporates support structure simulation, which includes estimating support volume, identifying risky faces, and assigning rewards based on overlap detection. This integrated approach facilitates the effective generation of 3D models with optimized support requirements.
@@ III Proposed Method
Our proposed method is built upon a pretrained diffusion-based 3D generative model $\pi_{\theta}(x|y)$, where $y$ is a text prompt and $x$ is the latent code for a mesh $\mathcal{M}$. Our method is to align $\pi_{\theta}$ with the preference of support-effective printing so that the generated mesh $\mathcal{M}$ requires minimal support structure volume $V(\mathcal{M}_{sup})$ when fabricated using common 3D printing processes.
@@ III-A Support Structure Simulation
In 3D printing, a model $\mathcal{M}$ is deposited layer by layer along a specified printing direction $\mathbf{d}$. Following the definitions of the maximal self-support angle $\alpha_{max}$ and self-supported regions as described in [3, 29, 4], we categorize the faces of $\mathcal{M}$ based on their printability. Specifically, for any face with normal vector $\mathbf{n}$, it is considered safe if it satisfies the following condition:
$\mathbf{n}\cdot\mathbf{d}+\sin{(\alpha_{max})}\geq 0,$ (1)
indicating that the face can be printed without additional support. Faces that do not meet this criterion are labeled as risky, as they are prone to requiring support structures during fabrication. To address this, segmentation-based methods have been developed to modify the geometry of $\mathcal{M}$, aiming to eliminate or reduce the presence of risky faces and thus improve overall printability. Although the risky angle depends on the printing device chosen, we use a widely-used value of $\alpha_{max}=45^{\circ}$ in the following context.
An illustrative example of risky face identification and support generation is presented in Fig. 2. The volumes of $\mathcal{M}$ and $\mathcal{M}_{sup}$ can be efficiently computed as $V(\mathcal{M})$ and $V(\mathcal{M}_{sup})$, respectively. However, since the volumes of different meshes can vary significantly, making fair comparisons of support effectiveness is challenging. To address this issue, we propose a simple yet effective metric called the normalized support volume (NSV), which normalizes the support volume by the volume of the mesh. Specifically, NSV is defined as:
@@
To efficiently evaluate NSV of support structures, we utilize Axis-Aligned Bounding Boxes (AABB) for rapid computation of the volume from risky faces to the print bed. This is achieved using the Embree [30], which facilitates fast AABB calculations to determine the spatial relationship between the model’s faces and the print surface. For each risky face, we compute its AABB and determine the intersection points with the print bed. If the risky face intersects with the print surface, we consider the volume above it as potential support volume. Given the small area of the face, we approximate the volume using three tetrahedra, which allows for a more accurate representation of the geometry. The volume $V$ of each tetrahedron can be computed using the formula:
$V=\frac{1}{6}\left|\begin{vmatrix}x_{1}&y_{1}&z_{1}&1\\ x_{2}&y_{2}&z_{2}&1\\ x_{3}&y_{3}&z_{3}&1\\ x_{4}&y_{4}&z_{4}&1\end{vmatrix}\right|$
where $(x_{i},y_{i},z_{i})$ are the coordinates of the tetrahedron’s vertices. The total volume $V_{\text{total}}$ from the three tetrahedra can then be expressed as:
$V_{\text{total}}=V_{1}+V_{2}+V_{3}$ (3)
By summing the volumes of the constructed tetrahedra, the simulation effectively estimates the total support volume extending from each risky face to either the print bed or its terminating height. This formulation allows the model to account for the structural stability required during fabrication. It is worth noting that our approach is agnostic to the specific support simulation method. In essence, the support simulation $S$ proceeds by (1) identifying risky faces, (2) locating their intersection points with either the print bed or the mesh surface, and (3) accumulating the volumes of all resulting tetrahedra.
@@ III-B Formulating Alignment with ODPO
We devise a solution to this problem by prompting the diffusion model to increase the likelihood of preferred responses compared to dispreferred ones through a penalization mechanism. This mechanism can be viewed as an ‘offset’ based on the difference in their respective reward values. To effectively optimize our model, we employ ODPO [6], which involves introducing an offset $\Delta o$ that accounts for the differences in the reward values between preferred and dispreferred samples. Specifically, we define $\Delta o$ based on the extent of preference indicated by the NSV values $r(y,x^{w})$ and $r(y,x^{l})$ for the preferred response $x^{w}$ and dispreferred response $x^{l}$, respectively.
@@
Figure 3: We present the printed results of models generated by the baseline approach (top), which necessitate substantial support structures during the printing process to ensure printability. In contrast, the models produced by our proposed SEG framework (bottom) demonstrate a significantly more efficient requirement for support structures.
@@ IV-A Experimental Setup
We train our proposed SEG model using $8\times$ NVIDIA A100 GPUs, each of which has 40GB VRAM, leveraging DeepSpeed [35] with ZeRO-1 offload for enhanced performance. The learning rate is configured to $5\times 10^{-6}$, and we utilize a batch size of 96 per iteration through gradient accumulation, completing a total of 50,000 training steps with the Adam optimizer [36]. Our foundation is the state-of-the-art 3D diffusion-based generative model TRELLIS (text-to-3D, XL size) [26], which serves as a robust starting point. To efficiently validate our proposed method, we use LoRA [37] training with rank of 64 and alpha of 128. For ODPO, the $\alpha$ is set to $1$. This approach enables effective fine-tuning while preserving the model’s original capabilities.
@@
To validate the effectiveness of our proposed method in terms of support-efficient 3D printing (3DP), we created two benchmark datasets. The first benchmark dataset, named Thingi10k-Val, is derived from the aforementioned curated dataset, where we split the entire Thingi10k [32] dataset into training and validation sets. The validation set consists of 100 prompts randomly selected from this split.
We also utilize GPT-4.1 to generate 60 prompts specifically designed for 3D generation in daily life, animation, and animal categories. These prompts were selected with a focus on scenarios likely to require support structures, ensuring that we account for cases that are too straightforward and do not necessitate supports, such as a simple cube. This dataset is referred to as GPT-3DP-Val in the following context.
@@
We utilize the NSV metric to evaluate the support structures generated by our proposed method and the baseline models, as it directly relates to the goal of support-efficient 3D generation. To extend NSV metric from a single model to the whole dataset, we define the volume-weighted NSV as the default computation method, and denote the arithmetic mean of NSV values as NSV. Additionally, different models may yield varying results for the same prompt, allowing us to assess the rate considering the cases of both win and tie (under a threshold of $10^{-3}$) based on NSV across the entire benchmark dataset, a measure we refer to as support-efficient consistency (SEC). This metric offers an alternative means of evaluating the consistency of support structure reduction while reducing the impact of randomness.
@@ IV-B Quantitative Results
[…]
TRELLIS 0.504 1.265 N/A
DPO 0.432 1.021 0.733
DRO 6.128 19.642 0.067
Ours 0.222 0.691 0.867
We conducted experiments on two validation datasets. The results of these evaluations are summarized in Tab. I and Tab. II. Our findings reveal that our method consistently outperforms all other evaluated techniques across both datasets. Notably, we observed a decline in performance following training with the DRO-based approach. We attribute this deterioration to the incompatibility of directly using NSV as a reward during training, as well as the absence of a reference model, which contributed to instability throughout the training process. These outcomes not only highlight the effectiveness and robustness of our proposed approach but also emphasize its potential to significantly advance the field of 3D generation.
@@
We present visualizations of the results generated by our method in comparison to the baseline model in Fig. 5. To enhance clarity, we have highlighted the areas deemed risky in red, indicating regions that necessitate the addition of support structures to prevent the collapse of overhanging sections during the printing process. Furthermore, we have fabricated several pairs of results using the Bambu Labs A1 3D printer, as depicted in Fig. 4. This practical demonstration, featured in Fig. 3, underscores the potential applications of our method in real-world physical printing scenarios. The ability to generate 3D models that maintain structural integrity while minimizing support requirements represents a significant advancement in the field of 3D model generation.

### v2

@@ Abstract
Current text-to-3D models prioritize visual fidelity but often neglect physical fabricability, resulting in geometries requiring excessive support structures. This paper introduces SEG (Support-Effective Generation), a novel framework that integrates Direct Preference Optimization with an Offset (ODPO) into the 3D generation pipeline to directly optimize models for minimal support material usage. By incorporating support structure simulation into the training process, SEG encourages the generation of geometries that inherently require fewer supports, thus reducing material waste and production time. We demonstrate SEG’s effectiveness through extensive experiments on two benchmark datasets, Thingi10k-Val and GPT-3DP-Val, showing that SEG significantly outperforms baseline models such as TRELLIS, DPO, and DRO in terms of support volume reduction and printability. Qualitative results further reveal that SEG maintains high fidelity to input prompts while minimizing the need for support structures. Our findings highlight the potential of SEG to transform 3D printing by directly optimizing models during the generative process, paving the way for more sustainable and efficient digital fabrication practices.
Index Terms:
Additive Manufacturing; AI and Machine Learning in Manufacturing and Logistics Systems; Manufacturing and Data Processing
@@ I Introduction
As generative models continue to mature, an increasing number of non-expert users can now create and fabricate 3D models directly from text prompts and accessible 3D printers, which becomes part of a broader democratization of personalized manufacturing. However, a significant ‘fabrication gap’ remains: current pipelines prioritize visual fidelity over physical manufacturability. A critical bottleneck is the generation of overhanging features that necessitate excessive support structures, leading to material waste, increased print time, and surface artifacts.
Existing text-to-3D models ignore printability constraints, while slicers handle support after geometry is fixed. These supports are crucial for maintaining the integrity of overhanging features during the printing process. Yet, they often lead to material waste and extended production times, and they can compromise the surface quality of the final product [5, 37]. Traditional geometric optimization approaches have proven effective by either deforming the 3D model in the post-processing phase [9] or decomposing the model into a sequence of subparts for multi-axis fabrication [34, 3]. Unfortunately, many existing 3D generation models fail to account for these practical constraints during their training, resulting in geometries that, while visually stunning, may not be feasible for real-world fabrication.
Recognizing these challenges, we set out to create a solution that not only enhances the capabilities of 3D generation but also aligns them with the practical realities of 3D printing. We name this type of pipeline as Semantic-Fabrication Co-Optimization, which allows the model to explore the manifold of semantically valid shapes to find one that also satisfies fabrication constraints. Specifically, we introduce SEG, a novel framework that employs Direct Preference Optimization with an Offset (ODPO) [1] to refine generative models specifically for support-efficient object creation. By integrating simulation-based feedback on support structure volume into the alignment objective, SEG encourages the generation of geometries that require fewer supports while maintaining the fidelity and aesthetic appeal of the designs. Our contributions can be summarized as follows:
@@ II-A Support-oriented Modeling and Optimization
Prior research in support-aware modeling primarily focused on post-processing strategies that modify a completed model to reduce the need for supports or to enhance its structural stability. For example, build-orientation optimization algorithms evaluate various possible orientations of a given part, seeking to minimize the area of unsupported faces or to reduce the average angle of overhangs [dumas2014efficient]. Once an orientation is selected, explicit support-generation methods design and position support structures—ranging from tree-like branches [26] to locally adaptive struts that are strategically placed according to geometric needs [27, 31]. These algorithms aim to uphold problematic features while balancing the trade-offs between print duration and material efficiency. Another class of techniques, partitioning-based methods, decompose complex models into smaller, more manageable sub-parts that can each be printed with fewer supports, subsequently reassembling them via geometric joints or adhesives [17].
By contrast, more recent approaches integrate support constraints directly into the modeling phase, rather than treating support generation as a secondary step. For instance, topology optimization under overhang constraints incorporates additional penalties—such as those based on feature angles or surface curvature—into classical density-based solvers, thereby discouraging the formation of unprintable regions during the design process [7]. Other methods operate directly on the mesh geometry, applying curvature-driven smoothing to eliminate sharp overhangs, or enforcing manufacturability criteria such as minimum wall thickness and maximum unsupported span [9].
Figure 1: The process begins with (a) data sampling from the Thingi10k database, where multi-view images are generated and captioned using Cap3D. The Text-to-Mesh transformation utilizes a sparse flow transformer (frozen) for model generation. (b) Subsequently, support structure simulation involves estimating the support volume, identifying risky faces, and assigning NSV rewards based on overlap detection. (c) The ODPO optimization leverages the NSV rewards to optimize the LoRAs (trainable) injected in the weighted-freeze sparse flow transformer.
@@ III Proposed Method
Our proposed method is built upon a pretrained diffusion-based 3D generative model $\pi_{\theta}(x|y)$, specifically TRELLIS [35], where $y$ is a text prompt and $x$ is the latent code for a mesh $\mathcal{M}$. The Sparse Flow Transformer refers to the backbone of the first generation stage in TRELLIS, which is a diffusion transformer variant modified for flow matching on sparse voxel tokens. It determines the topology and overall shape of the generated asset. TRELLIS operates in two distinct stages, both of which utilize transformer backbones: Stage 1: Sparse Structure Generation predicts the occupancy of a 3D voxel grid, which receives a text condition and predicts a sparse set of active voxels that define the coarse geometry of the object. It is “sparse” because it processes only the active voxel tokens rather than a dense volumetric grid, allowing for high resolution without cubic computational complexity. Stage 2: Structured Latent Decoding stage refines the features within the active voxels to generate detailed geometry and texture (appearance). Our method is to align $\pi_{\theta}$ with the preference of support-effective printing so that the generated mesh $\mathcal{M}$ requires minimal support structure volume $V(\mathcal{M}_{sup})$ when fabricated using common 3D printing processes.
@@ III-A Support Structure Simulation
In 3D printing, a model $\mathcal{M}$ is deposited layer by layer along a specified printing direction $\mathbf{d}$. Following the definitions of the maximal self-support angle $\alpha_{max}$ and self-supported regions as described in [9, 33, 34, 28, 11], we categorize the faces of $\mathcal{M}$ based on their printability. For every face $f$ in the mesh, calculate the dot product of its normal $\mathbf{n}$ with the gravity vector $\mathbf{g}=(0,0,-1)$. If $\mathbf{n}\cdot\mathbf{g}<\cos(135^{\circ})$ (assuming $45^{\circ}$ overhang threshold), the face is risky.
An illustrative example of risky face identification and support generation is presented in Fig. 2. The volumes of $\mathcal{M}$ and $\mathcal{M}_{sup}$ can be efficiently computed as $V(\mathcal{M})$ and $V(\mathcal{M}_{sup})$, respectively. However, since the volumes of different meshes can vary significantly, making fair comparisons of support effectiveness is challenging. To address this issue, we propose a simple yet effective metric called the normalized support volume (NSV). NSV serves as a scale-invariant proxy for the “printability cost”, which normalizes the support volume by the volume of the mesh. Specifically, NSV is defined as:
@@
From the centroid of each risky face, cast a ray in direction $\mathbf{g}$. We use a ray-tracing kernel [29] to find the first intersection point $P_{\text{int}}$. If $P_{\text{int}}$ is on the build plate ($z=0$), the support height is $z_{\text{face}}$. If $P_{\text{int}}$ is on the mesh itself (internal support), the support height is $z_{\text{face}}-z_{P_{\text{int}}}$. This per-face ray casting is robust to non-manifold geometry (holes, self-intersections) because it does not require a watertight solid to calculate volume.
The volume under a triangular face projected to a plane forms a triangular prism. Computing the volume of an arbitrary prism directly can be numerically unstable if the top and bottom bases are not perfectly parallel. Decomposing the prism into three tetrahedra is a standard method in Finite Element Method (FEM) [20] and computational geometry to calculate volumes of irregular polyhedra exactingly.
$V_{\text{prism}}=V_{\text{tet1}}+V_{\text{tet2}}+V_{\text{tet3}}$
where the volume of a tetrahedron with vertices $\mathbf{a},\mathbf{b},\mathbf{c},\mathbf{d}$ is $\frac{1}{6}|(\mathbf{a}-\mathbf{d})\cdot((\mathbf{b}-\mathbf{d})\times(\mathbf{c}-\mathbf{d}))|$.
It is worth noting that our approach is agnostic to the specific support simulation method. In essence, the support simulation $S$ proceeds by (1) identifying risky faces, (2) locating their intersection points with either the print bed or the mesh surface, and (3) accumulating the volumes of all resulting tetrahedra.
@@ III-B Formulating Alignment with ODPO
Standard DPO struggles when the preference margin between sample pairs is small or when the reward distribution is heavily skewed. To address this, we employ ODPO [1], which injects a dynamic margin $\Delta o$ into the loss function. This offset is proportional to the difference in NVS between the winning and losing samples. This explicitly penalizes the model more heavily when it fails to distinguish between a highly printable mesh and a support-heavy one, effectively scaling the gradient updates based on the magnitude of the geometric improvement. To effectively optimize our model, we employ ODPO [1], which involves introducing an offset $\Delta o$ that accounts for the differences in the reward values between preferred and dispreferred samples. Specifically, we define $\Delta o$ based on the extent of preference indicated by the NSV values $r(y,x^{w})$ and $r(y,x^{l})$ for the preferred response $x^{w}$ and dispreferred response $x^{l}$, respectively.
@@
Figure 3: Physical fabrication comparison using Fused Deposition Modeling (FDM). Top: Baseline (TRELLIS) outputs require dense support structures (blue regions in simulation) to support overhangs. Bottom: Our SEG framework generates self-supporting geometries that retain the input prompt’s semantics while minimizing material waste.
@@ IV-A Experimental Setup
Implementation Details.
In our SEG framework, the objective is to alter the geometry to reduce supports. Therefore, the LoRA adapters must be injected into the Sparse Flow Transformer. We train our model using $8\times$ NVIDIA A100 GPUs, each of which has 40GB VRAM, leveraging DeepSpeed [23] with ZeRO-1 offload for enhanced performance. The learning rate is configured to $5\times 10^{-6}$, and we utilize a batch size of 96 per iteration through gradient accumulation, completing a total of 50,000 training steps with the Adam optimizer [13]. Our foundation is the state-of-the-art 3D diffusion-based generative model TRELLIS (text-to-3D, XL size) [35], which serves as a robust starting point. To efficiently validate our proposed method, LoRA adapters [8] are attached to the Self-Attention and Cross-Attention layers of the transformer blocks. Specifically, we target the projection matrices for Query ($Q$), Key ($K$), Value ($V$), and Output ($O$). A rank of $r=64$ is typically sufficient for geometric adaptation in diffusion transformers, therefore, we specificy the rank as 64 and alpha as 128. For a pre-trained weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, the adapted weight is $W=W_{0}+\Delta W=W_{0}+BA$, where $B\in\mathbb{R}^{d\times r}$ and $A\in\mathbb{R}^{r\times k}$ are the trainable low-rank matrices. For ODPO, the $\alpha$ is set to $1$. This approach enables effective fine-tuning using $\mathcal{L}_{\text{ODPO}}$ (Eq. 4) while preserving the model’s original capabilities.
@@
To validate the effectiveness of our proposed method in terms of support-efficient 3D printing (3DP), we created two benchmark datasets. The first benchmark dataset, named Thingi10k-Val, is derived from the aforementioned curated dataset, where we split the entire Thingi10k [38] dataset into training and validation sets. The validation set consists of 100 prompts randomly selected from this split. We also utilize GPT-4.1 to generate 60 prompts specifically designed for 3D generation in daily life, animation, and animal categories. These prompts were selected with a focus on scenarios likely to require support structures, ensuring that we account for cases that are too straightforward and do not necessitate supports, such as a simple cube. This dataset is referred to as GPT-3DP-Val in the following context.
@@
We utilize the NSV metric to evaluate the support structures generated by our proposed method and the baseline models, as it directly relates to the goal of support-efficient 3D generation. To extend NSV metric from a single model to the whole dataset, we define the volume-weighted NSV as the default computation method, and denote the arithmetic mean of NSV values as NSV. Additionally, different models may yield varying results for the same prompt, allowing us to assess the rate considering the cases of both win and tie (under a threshold of $10^{-3}$) based on NSV across the entire benchmark dataset, a measure we refer to as support-efficient consistency (SEC). To evaluate the semantic fidelity, we computed CLIP scores [22] for the generated meshes.
@@ IV-B Quantitative Results
NSV $\downarrow$ NSV $^{*}\downarrow$ SEC $\uparrow$ CLIP $\uparrow$
TRELLIS 0.504 1.265 N/A 26.37
DPO 0.432 1.021 0.733 25.14
DRO 6.128 19.642 0.067 22.12
Ours 0.222 0.691 0.867 25.89
We conducted experiments on two validation datasets. The results of these evaluations are summarized in Tab. I and Tab. II. Our findings reveal that our method consistently outperforms all other evaluated techniques across both datasets. The DRO baseline exhibited a performance collapse. Unlike preference-based methods, which are anchored to a reference policy, DRO directly maximizes the reward. In the context of text-to-3D, this unconstrained optimization caused the model to drift into degenerate geometries that maximized volume or created unprintable artifacts, confirming that reference-guided alignment is essential for preserving semantic fidelity while optimizing for physics.
@@
We present visualizations of the results generated by our method in comparison to the baseline model in Fig. 5. To enhance clarity, we have highlighted the areas deemed risky in red, indicating regions that necessitate the addition of support structures to prevent the collapse of overhanging sections during the printing process. Furthermore, we have fabricated several pairs of results using the Bambu Labs A1 3D printer, as depicted in Fig. 4. This practical demonstration, featured in Fig. 3, underscores the potential applications of our method in real-world physical printing scenarios.
A key advantage of SEG is that it effectively learns by the NSV metric, which can measure relative differences, rather than requiring absolute precision. We validated this premise by performing a correlation analysis between our NSV metric and the support material volume calculated by the commercial slicer Cura. By comparing the weight ratios (with and without supports) of 50 models sampled from Thingi10k using Cura’s default ‘Normal’ settings, we observed a positive Pearson correlation coefficient ($r>0.85$), which confirms the effectiveness of our approximation.